# On endosplit $p$-permutation resolutions and Broué’s conjecture for $p$-solvable groups

### v2

@@ Abstract.
Endosplit $p$-permutation resolutions play an instrumental role in verifying Broué’s abelian defect group conjecture in numerous cases. In this article, we give a complete classification of endosplit $p$-permutation resolutions and reduce the question of Galois descent of an endosplit $p$-permutation resolution to the Galois descent of the module it resolves. This is shown using techniques from the study of endotrivial complexes, the invertible objects of the bounded homotopy category of $p$-permutation modules. As an application, we show that a refinement of Broué’s conjecture proposed by Kessar and Linckelmann in [KL18] holds for blocks of $p$-solvable groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$.
@@ 1. Introduction
If one can make a conjecture about blocks, Galois-invariance should be involved too. This philosophy motivates numerous recent refinements of the Alperin-McKay conjecture proposed by Issacs-Naavrro in [IN02], Navarro in [Nav04], and Turull in [Tur13], a refinement of Alperin’s weight conjecture proposed by Navarro in [Nav04], and a recent refinement of Broué’s abelian defect group conjecture proposed by Kessar-Linckelmann in [KL18]. The refinements of the Alperin-McKay conjecture add additional structure to the proposed bijections by imposing additional Galois-invariance or structural conditions. In particular, these refinements suggest that the conjecture should hold over non-splitting fields, descending to equivalences over finite fields of characteristic $p$ of any size. Kessar and Linckelmann’s refinement of the abelian defect group conjecture predicts exactly this: it predicts that the abelian defect group conjecture should hold over any choice of $p$-modular system, in particular, the $p$-modular system $(\mathbb{Q}_{p},\mathbb{Z}_{p},\mathbb{F}_{p}).$ Their refinement has been verified in numerous cases, including blocks with cyclic defect groups in [KL18], blocks with Klein-four defect groups in [Hua23], blocks of alternating groups in [Hua24], blocks of $\operatorname{SL}_{2}(q)$ and $\operatorname{GL}_{2}(q)$ in [HLZ23], and unipotent blocks of $\operatorname{GL}_{n}(q)$ in [HLZ24]. The refined conjecture also holds for blocks of symmetric groups, as the original proof provided in [CR08] holds independent of choice of field.
The goal of this paper is twofold. Our first objective is to provide additional evidence that Kessar and Linckelmann’s refinement of Broué’s conjecture is true by showing it holds for blocks of a $p$-solvable groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$, the key reduction step in Harris and Linckelmann’s verification of the conjecture for all blocks of $p$-solvable groups [HL00]. This is a small improvement in the literature, a step up from the result of [Mil25b] in which we verified the conjecture for blocks of $p$-nilpotent groups with abelian Sylow $p$-subgroups. Many of the prior verifications of the refined conjecture are shown by taking a previous construction of a splendid Rickard equivalence and demonstrating that it satisfies a Galois descent condition. We carry out the same process here, taking the explicit construction in [HL00] and demonstrating that it descends. Unfortunately, we are not able to show that the reduction to the $G=O_{p^{\prime},p,p^{\prime}}(G)$ case also satisfies Galois descent; we remark on the key obstruction to our strategy at this at the end of the paper.
@@
(Theorem 5.12) Let $G$ be a finite $p$-solvable group satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$, and suppose $b$ be a block idempotent of $\mathbb{F}_{p}G$ with abelian defect group. Then $b$ is splendidly Rickard equivalent to its Brauer correspondent $c$. In particular, Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture holds for blocks of $p$-solvable groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$.
Harris and Linckelmann’s construction relies on the well-understood block theory of $p$-solvable groups, and uses the machinery of endosplit $p$-permutation resolutions. Such resolutions were introduced by Rickard in [Ric96] in his verification of the abelian defect group conjecture for $p$-nilpotent groups, and have since become objects of independent interest. The second objective, and perhaps the main focus of this paper, is to give a complete characterization of all endosplit $p$-permutation resolutions.
@@
[Ric96] Let $G$ be a finite group. An endosplit $p$-permutation resolution of a finitely generated $kG$-module $M$ is a bounded complex $C$ of $p$-permutation $kG$-modules together with an isomorphism between $M$ and the degree zero homology of $C$ such that
@@
    the homology of $M$ is concentrated in degree zero, and
@@
Moreover, we prove that the question of Galois descent of endosplit $p$-permutation resolutions reduces to the question of Galois descent of the modules they resolve. If $k^{\prime}/k$ is a field extension, we say a $k^{\prime}G$-module $M^{\prime}$ descends to $k$ if there exists a $kG$-module $M$ such that $k^{\prime}\otimes_{k}M\cong M^{\prime}$, and similarly for chain complexes. Tensor products of relative syzygies are defined at Definition 3.15 - they can be thought of “invertible” endosplit $p$-permutation resolutions which exist independent of field, and have an associated group that parametrizes them, which we describe. We remark that normally, relative syzygies refer to modules that arise as the nonzero homology of chain complexes. In this paper, by a relative syzygy, we mean the corresponding chain complex.
Our description is heavily inspired by Dade’s description of all endopermutation $kP$-modules for finite $p$-groups $P$ in terms of the “capped” endopermutation $kP$-modules [Dad78]. In the module-theoretic case, the question of classifying the capped endopermutation modules is a thoroughly challenging task, and was closed out nearly 30 years later by Bouc [Bou06]. However in our case, tensor products of relative syzygies are easily described as products of certain two-term chain complexes.
@@
    Let $C$ be an endosplit $p$-permutation resolution of $kG$-modules. Then $C$ is a direct summand of the endosplit $p$-permutation resolution $\bigoplus_{i=1}^{l}\operatorname{Ind}^{G}_{P_{i}}\operatorname{Res}^{G}_{P_{i}}Y$ for some tensor product of relative syzygies $Y$ and family of $p$-subgroups $\{P_{i}\}_{i=1}^{l}$ of $G$.
@@
    Let $k^{\prime}/k$ be an extension of finite fields. Let $C$ be an endosplit $p$-permutation resolution for an endo-$p$-permutation $k^{\prime}G$-module $M$ which contains no non-zero contractible summands. Then $C$ descends to $k$ if and only if $M$ does.
The paper is organized as follows. Section 2 recalls some prerequisite information on projectivity relative to a module and relatively endotrivial modules. Section 3 concerns endosplit $p$-permutation resolutions and relatively endotrivial complexes, and builds up to the classification of endosplit $p$-permutation resolutions. Section 4 concerns the Galois descent of endosplit $p$-permutation resolutions. Finally, Section 5 applies the results of the previous sections towards verifying Kessar and Linckelmann’s refinement of Broué’s conjecture.
Notation and conventions: Throughout this paper, $G$ will be a finite group, $p$ a prime, and $(K,\mathcal{O},k)$ will denote a $p$-modular system. That is, $\mathcal{O}$ is a complete discrete valuation ring, $k$ is its residue field, and $K$ is its field of fractions. Denote by $s_{p}(G)$ the set of $p$-subgroups of $G$, and $\operatorname{Syl}_{p}(G)$ the set of Sylow $p$-subgroups of $G$. The additive group of class functions valued on $p$-subgroups $s_{p}(G)\to\mathbb{Z}$ is denoted $CF(G,p)$.
All modules and algebras are finitely generated, and all chain complexes are bounded. We denote the category of finitely generated $kG$-modules by ${}_{kG}\mathbf{mod}$, and the category of finitely generated $p$-permutation $kG$-modules by ${}_{kG}\mathbf{triv}$. Given an additive category $\mathcal{A}$, $Ch^{b}(\mathcal{A})$ denotes the bounded chain complex category of $\mathcal{A}$, and $K^{b}(\mathcal{A})$ denotes the homotopy category of $\mathcal{A}$. The symbol $\simeq$ denotes homotopy equivalence of chain complexes, and given a $kG$-module $M$, we write $M[i]$ to denote the chain complex with $M$ in degree $i$ and zero in all other degrees.
Given any $P$-subgroup $P$ of $G$, the Brauer construction is denoted as a functor $-(P):{}_{kG}\mathbf{triv}\to{}_{k[N_{G}(P)/P]}\mathbf{triv}$ or $Ch^{b}({}_{kG}\mathbf{triv})\to Ch^{b}({}_{k[N_{G}(P)/P]}\mathbf{triv})$. The Brauer homomorphism is denoted $\operatorname{Br}_{P}:(kG)^{P}\to kC_{G}(P)$.
For this paper, we assume familiarity with splendid Rickard equivalences, $p$-permutation modules, the Brauer construction, and some standard block theory and homological algebra; we refer the reader to [Lin18a] and [Lin18b] for a detailed treatment of these topics. We refer the reader to [Bou10, Section 12] for an overview of endopermutation modules for $p$-groups and the Dade group.
@@ 2. Relative $V$-projectivity and relatively endotrivial modules
For this section, $V$ is a $kG$-module, possibly 0. We review the notion of projectivity relative to a module, which was first considered by Okuyama in an unpublished manuscript [Oku], and relative endotriviality, which was first introduced by Lassueur in [Las11]. We also introduce the generalized Dade group, introduced by Lassueur in [Las13], which will be relevant in the sequel.
@@
    $M$ is relatively $V$-projective if there exists a $kG$-module $N$ for which $M\mid V\otimes_{k}N$. Denote the full subcategory of ${}_{kG}\mathbf{mod}$ consisting of $V$-projective modules by ${}_{kG}\mathbf{mod}(V),$ and denote by ${}_{kG}\mathbf{triv}(V)$ the full subcategory consisting of $p$-permutation $V$-projective modules.
@@
    [Las11, Definition 3.1.1] $M$ is relatively $V$-endotrivial if $M^{*}\otimes_{k}M\cong k\oplus N$ for some relatively $V$-projective $kG$-module $N$.
@@
    [Las13, Definition 5.3] Set $V(\mathcal{F}_{G}):=\bigoplus_{P\in s_{p}(G)\setminus\operatorname{Syl}_{p}(G)}k[G/P]$. $M$ is a strongly capped endo-$p$-permutation $kG$-module if $M$ is endo-$p$-permutation, i.e. $M^{*}\otimes_{k}M$ is $p$-permutation, and $M$ is $V(\mathcal{F}_{G})$-endotrivial. If $G$ is a $p$-group, one may easily verify $M$ is strongly capped endo-$p$-permutation if and only if $M$ is capped endopermutation, i.e. $M^{*}\otimes_{k}M$ is a permutation module with the trivial module as a direct summand.
@@
Conversely, if $V$ is $p$-permutation, let $\mathcal{X}_{V}$ denote the set of all $P\in s_{p}(G)$ for which $V(P)\neq 0$. Then projectivity relative to $V$ is equivalent to projectivity relative to $\mathcal{X}_{V}$; see [Mil25c, Theorem 3.7(c)]. In particular, if $\mathcal{X}\subset s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$, then $\mathcal{X}$-projectivity implies $V(\mathcal{F}_{G})$-projectivity.
@@
[Las11, Proposition 2.2.2] Let $k$ be an algebraically closed field. The following are equivalent.
@@
    ${}_{kG}\mathbf{mod}(V)\subset{}_{kG}\mathbf{mod}$.
@@
For this reason, relative $V$-projectivity is interesting only when $V$ is absolutely $p$-divisible. We remark that in the previous proposition, $k$ is assumed to be algebraically closed in [Las11]. However, this will not pose a problem in the sequel, as we will primarily on relative projectivity with respect to the module $V(\mathcal{F}_{G})$, and it is straightforward to verify that all three of the above conditions hold for this module, regardless of algebraic closure.
@@
[Las11, Lemma 3.3.1] Let $M$ be a relatively $V$-endotrivial $kG$-module with $V$ absolutely $p$-divisible. $M\cong M_{0}\oplus U$, where $M_{0}$ is an indecomposable relatively $V$-endotrivial $kG$-module and $U$ is a $V$-projective $kG$-module or 0. Moreover, $M_{0}$ has vertex set $\operatorname{Syl}_{p}(G)$.
@@
Let $V$ be an absolutely $p$-divisible $kG$-module.
@@
    [Las11, Proposition 3.5.1] We impose an equivalence relation $\sim_{V}$ on the class of $V$-endotrivial $kG$-modules as follows: if $M$ and $N$ are $V$-endotrivial, write $M\sim_{V}N$ if and only if $M$ and $N$ have isomorphic indecomposable $V$-endotrivial summands. Let $T_{k,V}(G)$ denote the resulting set of equivalence classes. $T_{k,V}(G)$ is an abelian group with addition defined by $[M]+[N]:=[M\otimes_{k}N]$. We have $-[M]=[M^{*}]$ and each equivalence class has a unique indecomposable representative.
@@
    [Las13, Definition 5.5] Let $D_{k}(G)$ denote the set of equivalence classes of strongly capped endo-$p$-permutation $kG$-modules under the equivalence relation $\sim_{V(\mathcal{F}_{G})}$. As before, $D_{k}(G)$ forms an abelian group with addition defined by $[M]+[N]:=[M\otimes_{k}N]$. Moreover, $D_{k}(G)$ identifies as a subgroup of $T_{V(\mathcal{F}_{G})}(G)$. $D_{k}(G)$ is the generalized Dade group of $G$. If $G$ is a $p$-group, $D_{k}(G)$ we recover the classical Dade group $D_{k}(G)$. We refer the reader to [Bou10, Section 12] for a comprehensive treatment of the classical Dade group.
@@
    Given any $G$-set $X$, write $\Delta(X)$ for the $kG$-module given by the kernel of the augmentation homomorphism $kX\to k$. Let $S\in\operatorname{Syl}_{p}(G)$. If $X^{S}=0$, then $\Delta(X)$ is an endo-$p$-permutation module, see [Alp01]. We write $\Omega_{X}$ for the equivalent class of $\Delta(X)$ in $D_{k}(G)$.
    Extending notation, write $\Delta(X)^{-1}$ for the $kG$-module given by the cokernel of the inclusion $k\to kX$. We have $[\Delta(X)^{-1}]=-\Omega_{X}\in D_{k}(G)$. We define $D^{\Omega}(G)$ to be the subgroup of $D_{k}(G)$ generated by elements of the form $\Omega_{X}$, where $X$ runs over all $G$-sets $X$ satisfying $X^{S}=0$. In fact, $D^{\Omega}(G)$ is generated by elements of the form $\Omega_{G/P}$, where $P$ runs over a set of conjugacy class of representatives of $s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$, see [Las13, Lemma 12.1]. Note that $D^{\Omega}(G)$ does not depend on the size of the base field $k$. For $P\in s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$, we write $\Omega_{P}:=\Omega_{G/P}$ for shorthand.
@@
[Mil25c, Theorem 3.7(c)] Let $V$ be a $kG$-module. Then, there exists a $p$-permutation module $W$ such that ${}_{kG}\mathbf{triv}(V)={}_{kG}\mathbf{triv}(W)$. In particular, there exists a poset of $p$-subgroups $\mathcal{X}\subseteq s_{p}(G)$ closed under $G$-conjugation and taking subgroups for which ${}_{kG}\mathbf{triv}(V)={}_{kG}\mathbf{triv}(\mathcal{X})$.
@@
First we show that if $M$ is a trivial source module with vertex $P\in s_{p}(G)$ which belongs to ${}_{kG}\mathbf{triv}(V)$, then $k[G/P]$ belongs to ${}_{kG}\mathbf{triv}(V)$. If $M\mid V\otimes_{k}N$ for some $kG$-module $N$, then $\operatorname{Res}^{G}_{S}M\mid\operatorname{Res}^{G}_{S}(V\otimes_{k}N)$ for $S\in\operatorname{Syl}_{p}(G)$. Since $M$ has vertex $P$, $\operatorname{Res}^{G}_{S}M$ has a direct summand isomorphic to $k[S/P]$, and any other nonisomorphic direct summand is isomorphic to $k[S/Q]$ for some subgroup $Q$ $G$-conjugate to a subgroup of $P$. Therefore $\operatorname{Res}^{G}_{S}(V\otimes_{k}N)$ has a direct summand isomorphic to $k[S/P]$. We have $\operatorname{Res}^{G}_{S}(V\otimes_{k}N)\cong k[S/P]\oplus N^{\prime}$, for some $kS$-module $N$, therefore, $k[S/P]$ is $\operatorname{Res}^{G}_{S}V$-projective. Therefore by [Las12, Proposition 2.2.2], ${}_{kS}\mathbf{triv}(P)\subseteq{}_{kS}\mathbf{triv}(\operatorname{Res}^{G}_{S}V)$. This holds for all $G$-conjugates of $P$.
@@
Now, let $\mathcal{X}$ be the set of all vertices which occur for all trivial source modules belonging to ${}_{kG}\mathbf{triv}(V)$. The previous claim and [Las12, Proposition 2.2.2] imply that
@@
The converse inclusion follows from [Mil25c, Proposition 3.5]. ∎

### v3

@@ Abstract.
Endosplit $p$-permutation resolutions play an instrumental role in verifying Broué’s abelian defect group conjecture in numerous cases. We give a new characterization of all endosplit $p$-permutation resolutions and reduce the question of Galois descent of an endosplit $p$-permutation resolution to the Galois descent of the module it resolves. This is shown using techniques from the study of endotrivial complexes, the invertible objects of the bounded homotopy category of $p$-permutation modules. As an application, we show that a refinement of Broué’s conjecture proposed by Kessar–Linckelmann holds for certain blocks of groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with abelian Sylow $p$-subgroup, the key reduction step in Harris–Linckelmann’s verification of Broué’s conjecture for all $p$-solvable groups.
@@ 1. Introduction
If one can make a conjecture about blocks, Galois-invariance should be involved too. This philosophy motivates numerous recent refinements of the Alperin–McKay conjecture proposed by Isaacs–Navarro in [IN02], Navarro in [N04], and Turull in [T13], a refinement of Alperin’s weight conjecture proposed by Navarro in [N04], and a recent refinement of Broué’s abelian defect group conjecture proposed by Kessar–Linckelmann in [KL18]. The refinements of the Alperin–McKay conjecture add additional structure to the proposed bijections by imposing additional Galois-invariance or structural conditions. In particular, these refinements suggest that the conjecture should hold over non-splitting fields, descending to equivalences over finite fields of characteristic $p$ of any size. Kessar–Linckelmann’s refinement of the abelian defect group conjecture predicts exactly this: it predicts that the abelian defect group conjecture should hold over any choice of $p$-modular system, in particular, the $p$-modular system $(\mathbb{Q}_{p},\mathbb{Z}_{p},\mathbb{F}_{p}).$ Their refinement has been verified in numerous cases, including blocks with cyclic defect groups in [KL18], blocks with Klein-four defect groups in [H23], blocks of alternating groups in [H24c], blocks of $\operatorname{SL}_{2}(q)$ and $\operatorname{GL}_{2}(q)$ in [HLZ23], and unipotent blocks of $\operatorname{GL}_{n}(q)$ in [HLZ23b]. The refined conjecture also holds for blocks of symmetric groups, as the original proof provided in [CR08] holds independent of choice of field.
The goal of this paper is twofold. Our first minor objective is to provide additional evidence that Kessar–Linckelmann’s refinement of Broué’s conjecture is true by showing it holds for particular blocks of $p$-solvable groups $G$, satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$, with abelian Sylow $p$-subgroup. The case we consider is the key reduction step in Harris–Linckelmann’s verification of the conjecture for all blocks of $p$-solvable groups [HL00]. This is a minor improvement in the literature, a step up from the result of [M24d] in which we verified the conjecture for blocks of $p$-nilpotent groups with abelian Sylow $p$-subgroups. However, if one can prove an analogue of [HL00, Proposition 3.1] over arbitrary fields of characteristic $p$, a verification of Kessar–Linckelmann’s conjecture for all $p$-solvable groups could follow.
@@
(Theorem 5.12) Let $G$ be a $p$-solvable group satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with a abelian Sylow $p$-subgroup $P$. Let $b^{\prime}$ be a $G$-stable block of $\overline{\mathbb{F}}_{p}O_{p^{\prime}}(G)$, and let $b$ be a block idempotent of $\mathbb{F}_{p}G$ for which $bb^{\prime}\neq 0$. Let $c$ be the Brauer correspondent of $b$ in $\mathbb{F}_{p}N_{G}(P)$. Then $b$ and $c$ are splendidly Rickard equivalent.
Harris–Linckelmann’s construction relies on the well-understood block theory of $p$-solvable groups, and uses the machinery of endosplit $p$-permutation resolutions.
@@
[R96] Let $G$ be a finite group. An endosplit $p$-permutation resolution of a finitely generated $kG$-module $M$ is a bounded complex $C$ of $p$-permutation kGkG-modules together with an isomorphism between $M$ and the degree zero homology of $C$ such that
@@
    the homology of $C$ is concentrated in degree zero, and
@@
Such resolutions were introduced by Rickard in his verification of the abelian defect group conjecture for $p$-nilpotent groups, and have since become objects of independent interest. Our second objective, and main achievement of this paper, is to give a complete characterization of all endosplit $p$-permutation resolutions. Moreover, we prove that the question of Galois descent of endosplit $p$-permutation resolutions reduces to the question of Galois descent of the modules they resolve. If $k^{\prime}/k$ is a field extension, we say a $k^{\prime}G$-module $M^{\prime}$ descends to $k$ if there exists a $kG$-module $M$ such that $k^{\prime}\otimes_{k}M\cong M^{\prime}$, and similarly for chain complexes.
@@
    Let $C$ be an endosplit $p$-permutation resolution with contractible no direct summands of a $kG$-module. Then $C$ is a direct summand of the endosplit $p$-permutation resolution $\bigoplus_{i=1}^{l}\operatorname{Ind}^{G}_{P_{i}}\operatorname{Res}^{G}_{P_{i}}Y$ for some chain tensor of relative syzygies $Y$ (see Definition 3.15) and family of $p$-subgroups $\{P_{i}\}_{i=1}^{l}$ of $G$.
@@
We note that normally, relative syzygies refer to modules that arise as the nonzero homology of chain complexes. In this paper, by a relative syzygy, we mean the corresponding chain complex. Our description is heavily inspired by Dade’s description of all endopermutation $kP$-modules for finite $p$-groups $P$ in terms of the capped endopermutation $kP$-modules [D78a]. In the module-theoretic case, the question of classifying the capped endopermutation modules is a thoroughly challenging task, and was closed out nearly 30 years later by Bouc [Bo06]. However in our case, tensor products of relative syzygies are easily described as products of certain two-term chain complexes.
The paper is organized as follows. Section 2 recalls some prerequisite information on projectivity relative to a module and relatively endotrivial modules. Section 3 concerns endosplit $p$-permutation resolutions and relatively endotrivial complexes, and builds up to the classification of endosplit $p$-permutation resolutions. Section 4 concerns the Galois descent of endosplit $p$-permutation resolutions. Finally, Section 5 applies the results of the previous sections towards verifying Kessar–Linckelmann’s refinement of Broué’s conjecture.
Notation and conventions: Throughout this paper, $G$ will be a finite group, $p$ a prime, and $(K,\mathcal{O},k)$ will denote a $p$-modular system. That is, $\mathcal{O}$ is a complete discrete valuation ring, $k$ is its residue field, and $K$ is its field of fractions. Denote by $s_{p}(G)$ the set of $p$-subgroups of $G$, and $\operatorname{Syl}_{p}(G)$ the set of Sylow $p$-subgroups of $G$. The additive group of superclass functions valued on $p$-subgroups $s_{p}(G)\to\mathbb{Z}$ is denoted $CF(G,p)$.
All modules and algebras are assumed finitely generated, and all chain complexes are assumed bounded. We denote the category of finite $G$-sets by $_{G}\mathbf{set}$, the category of finitely generated $kG$-modules by $_{kG}\mathbf{mod}$, and the category of finitely generated $p$-permutation kGkG-modules by $_{kG}\mathbf{triv}$. Given an additive category $\mathcal{A}$, $Ch^{b}(\mathcal{A})$ denotes the bounded chain complex category of $\mathcal{A}$, and $K^{b}(\mathcal{A})$ denotes the bounded homotopy category of $\mathcal{A}$. The symbol $\simeq$ denotes homotopy equivalence of chain complexes, i.e., isomorphism in $K^{b}(\mathcal{A})$, and given a $kG$-module $M$, we write $M[i]$ to denote the chain complex with $M$ in degree $i$ and zero in all other degrees.
[…]
For this paper, we assume familiarity with splendid Rickard equivalences, $p$-permutation modules, the Brauer construction, and some standard block theory and homological algebra; we refer the reader to [L181] and [L182] for a detailed treatment of these topics. We refer the reader to [Bou10, Section 12] for an overview of endopermutation modules for $p$-groups and the Dade group.
@@ 2. Relative $V$-projectivity and relatively endotrivial modules
For this section, $V$ is a $kG$-module, possibly 0. We review the notion of projectivity relative to a module, which was first considered by Okuyama in an unpublished manuscript [O], and relative endotriviality, which was first introduced by Lassueur in [CL11]. We also introduce the generalized Dade group, introduced by Lassueur in [CL13], which will be relevant in the sequel.
@@
    $M$ is relatively $V$-projective if there exists a $kG$-module NN for which $M\mid V\otimes_{k}N$. Denote the full subcategory of $_{kG}\mathbf{mod}$ consisting of $V$-projective modules by $_{kG}\mathbf{mod}(V),$ and denote by $_{kG}\mathbf{triv}(V)$ the full subcategory consisting of $p$-permutation $V$-projective modules.
@@
    [CL11, Definition 3.1.1] $M$ is relatively $V$-endotrivial if $M^{*}\otimes_{k}M\cong k\oplus N$ for some relatively $V$-projective kGkG-module NN.
@@
    [CL13, Definition 5.3] Set $V(\mathcal{F}_{G}):=\bigoplus_{P\in s_{p}(G)\setminus\operatorname{Syl}_{p}(G)}k[G/P]$. $M$ is a strongly capped endo-$p$-permutation kGkG-module if $M$ is endo-$p$-permutation, i.e. $M^{*}\otimes_{k}M$ is $p$-permutation, and $M$ is $V(\mathcal{F}_{G})$-endotrivial. If $G$ is a $p$-group, one may easily verify $M$ is strongly capped endo-$p$-permutation if and only if $M$ is capped endopermutation, i.e. $M^{*}\otimes_{k}M$ is a permutation module with the trivial module as a direct summand.
@@
Conversely, if $V$ is $p$-permutation, let $\mathcal{X}_{V}$ denote the set of all $P\in s_{p}(G)$ for which $V(P)\neq 0$. Then projectivity relative to $V$ is equivalent to projectivity relative to $\mathcal{X}_{V}$; see [M24b, Theorem 3.7(c)]. In particular, if $\mathcal{X}\subseteq s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$, then $\mathcal{X}$-projectivity implies $V(\mathcal{F}_{G})$-projectivity.
@@
[CL11, Propositions 2.2.2 and 3.4.1(a)] Let $k$ be an algebraically closed field. The following are equivalent.
@@
    $_{kG}\mathbf{mod}(V)\subsetneq{}_{kG}\mathbf{mod}$.
@@
For this reason, relative $V$-projectivity is interesting only when $V$ is absolutely $p$-divisible. We remark that in the previous proposition, $k$ is assumed to be algebraically closed in [CL11]. However, this will not pose a problem in the sequel, as we will primarily on relative projectivity with respect to the module $V(\mathcal{F}_{G})$, and it is straightforward to verify that all three of the above conditions hold for this module, regardless of algebraic closure.
@@
[CL11, Lemma 3.3.1] Let $M$ be a relatively $V$-endotrivial kGkG-module with $V$ absolutely $p$-divisible. $M\cong M_{0}\oplus U$, where $M_{0}$ is an indecomposable relatively $V$-endotrivial kGkG-module and $U$ is a $V$-projective kGkG-module or 0. Moreover, $M_{0}$ has vertex set $\operatorname{Syl}_{p}(G)$.
@@
Let $V$ be an absolutely $p$-divisible kGkG-module.
@@
    [CL11, Proposition 3.5.1] We impose an equivalence relation $\sim_{V}$ on the class of $V$-endotrivial kGkG-modules as follows: if $M$ and $N$ are $V$-endotrivial, write $M\sim_{V}N$ if and only if $M$ and $N$ have isomorphic indecomposable $V$-endotrivial summands. Let $T_{k,V}(G)$ denote the resulting set of equivalence classes. $T_{k,V}(G)$ is an abelian group with addition defined by $M]+[N]:=[M\otimes_{k}N]$. We have $M]=[M^{*}]$ and each equivalence class has a unique indecomposable representative.
@@
    [CL13, Definition 5.5] Let $D_{k}(G)$ denote the set of equivalence classes of strongly capped endo-$p$-permutation kGkG-modules under the equivalence relation $\sim_{V(\mathcal{F}_{G})}$. As before, $D_{k}(G)$ forms an abelian group with addition defined by $M]+[N]:=[M\otimes_{k}N]$. Moreover, $D_{k}(G)$ identifies as a subgroup of $T_{k,V(\mathcal{F}_{G})}(G)$. $D_{k}(G)$ is the generalized Dade group of $G$. If $G$ is a $p$-group, $D_{k}(G)$ we recover the classical Dade group $D_{k}(G)$. We refer the reader to [Bou10, Section 12] for a comprehensive treatment of the classical Dade group.
@@
    Given any $G$-set $X$, write $\Delta(X)$ for the $kG$-module given by the kernel of the augmentation homomorphism $kX\to k$. Let $S\in\operatorname{Syl}_{p}(G)$. If $X^{S}\neq\emptyset$, then $\Delta(X)$ is an endo-$p$-permutation module, see [Al01]. We write $\Omega_{X}$ for the equivalent class of $\Delta(X)$ in $D_{k}(G)$.
    Extending notation, write $\Delta(X)^{-1}$ for the $kG$-module given by the cokernel of the inclusion $k\to kX$. We have $\Delta(X)^{-1}]=-\Omega_{X}\in D_{k}(G)$. We define $D^{\Omega}(G)$ to be the subgroup of $D_{k}(G)$ generated by elements of the form $\Omega_{X}$, where $X$ runs over all $G$-sets $X$ satisfying $X^{S}=\emptyset$. In fact, $D^{\Omega}(G)$ is generated by elements of the form $\Omega_{G/P}$, where $P$ runs over a set of conjugacy class of representatives of $s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$, see [CL13, Lemma 12.1]. Note that $D^{\Omega}(G)$ does not depend on the size of the base field $k$. For $P\in s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$, we write $\Omega_{P}:=\Omega_{G/P}$ for shorthand.
@@
[M24b, Theorem 3.7(c)] Let $V$ be a $kG$-module. Then, there exists a $p$-permutation module $W$ such that $_{kG}\mathbf{triv}(V)={}_{kG}\mathbf{triv}(W)$. In particular, there exists a poset of $p$-subgroups $\mathcal{X}\subseteq s_{p}(G)$ closed under $G$-conjugation and taking subgroups for which $_{kG}\mathbf{triv}(V)={}_{kG}\mathbf{triv}(\mathcal{X})$.
@@
First we show that if $M$ is a trivial source module with vertex $P\in s_{p}(G)$ which belongs to $_{kG}\mathbf{triv}(V)$, then $k[G/P]$ belongs to $_{kG}\mathbf{triv}(V)$. If $M\mid V\otimes_{k}N$ for some $kG$-module NN, then $\operatorname{Res}^{G}_{S}M\mid\operatorname{Res}^{G}_{S}(V\otimes_{k}N)$ for $S\in\operatorname{Syl}_{p}(G)$. Since $M$ has vertex $P$, $\operatorname{Res}^{G}_{S}M$ has a direct summand isomorphic to $k[S/P]$, and any other nonisomorphic direct summand is isomorphic to $k[S/Q]$ for some subgroup $Q$ $G$-conjugate to a subgroup of $P$. Therefore $\operatorname{Res}^{G}_{S}(V\otimes_{k}N)$ has a direct summand isomorphic to $k[S/P]$. We have $\operatorname{Res}^{G}_{S}(V\otimes_{k}N)\cong k[S/P]\oplus N^{\prime}$, for some $kS$-module NN, therefore, $k[S/P]$ is $\operatorname{Res}^{G}_{S}V$-projective. Therefore by [CL12, Proposition 2.2.2], $_{kS}\mathbf{triv}(P)\subseteq{}_{kS}\mathbf{triv}(\operatorname{Res}^{G}_{S}V)$. This holds for all $G$-conjugates of $P$.
@@
Now, let $\mathcal{X}$ be the set of all vertices which occur for all trivial source modules belonging to $_{kG}\mathbf{triv}(V)$. The previous claim and [CL12, Proposition 2.2.2] imply that
@@
The converse inclusion follows from [M24b, Proposition 3.5]. ∎